If sin theta + cos theta=a then sin^4 theta + cos^4 theta=

Grade:10

2 Answers

Anushka Mishra
6 years ago
sin theta + cos theta=a(sin^2 theta)^2 + (cos^2 theta)^2=(sin^2 theta + cos^2 theta)^2 - 2sin^2 theta cos^2 theta1 - 2[ (sin theta + cos theta)^2 - (sin^2 theta + cos^2 theta)/2]^2= 1- 2[( a^2-1)^2/4]= 1- (a^2-1)^2/2You can expand it now.
Karthikeya I C4IPL
2 years ago
sin theta + cos theta=a(sin^2 theta)^2 + (cos^2 theta)^2=(sin^2 theta + cos^2 theta)^2 - 2sin^2 theta cos^2 theta1 - 2[ (sin theta + cos theta)^2 - (sin^2 theta + cos^2 theta)/2]^2= 1- 2[( a^2-1)^2/4]= 1- (a^2-1)^2/2You can expand it now.
This is the answer